OpenStudy (anonymous):

how many three digit number can be formed from 1,2,3,4 and 5 if repetition is allowed?

OpenStudy (anonymous):

sorry brother sister hard for me to understand english + math

OpenStudy (raden):

i cant speak english (i from indonesia), but i love MATH :p

OpenStudy (anonymous):

your good me nosebleed sometime hahaha

OpenStudy (anonymous):

5^3

OpenStudy (anonymous):

rad, can you help with parametric equations

OpenStudy (anonymous):

thanks eddie

OpenStudy (raden):

right 5^3 = 125 what ur question eddie

OpenStudy (anonymous):

i want to convert a parametric equation to a polar

OpenStudy (raden):

for example ?

OpenStudy (anonymous):

, x = 4sin t + 2sin(2t) , y = 4cos (t) + 2cos(2t)

OpenStudy (anonymous):

i tried doing x^2 + y^2 , but it wouldn't eliminate t. the directions say change the parametric equation to polar equation

OpenStudy (anonymous):

how many four digits number can be formed from 1,2,3,4 and 5 if no repetition is allowed?

OpenStudy (anonymous):

5 choices for first number, then 4 choices, then 3 choices, then 2 choices 5x4x3x2

OpenStudy (anonymous):

thanks eddie

OpenStudy (raden):

coordinat polar like (r, alpha) ? with r=sqrt(x^2+y^2) alpha = arctan (y/x), right eddie ?

OpenStudy (anonymous):

, yes

OpenStudy (anonymous):

so my idea was to change parametric to cartesian, then cartesian to polar. but it didnt work

OpenStudy (anonymous):

if there are three routes from manila to pampanga and also three from pampanga to pangasinan.How many different routes may be taken if one wants to travel from manila to pangasinan?

OpenStudy (anonymous):

3x3

OpenStudy (anonymous):

you finish my assigment eddie 3 more to go yehey =))))

OpenStudy (anonymous):

wahahaha

OpenStudy (raden):

x^2 = [4sin t + 2sin(2t)]^2 = 16sin^2 t + 16sintsin2t + 4sin^2 2t y^2 = [4cos (t) + 2cos(2t)]^2 = 16cos^2 t + 16costcos2t + 4cos^2 2t x^2+y^2 = 16(sin^2 t + cos^2 t) + 16(sintsin2t+costcos2t) + 4(sin^2 2t + cos^2 2t) = 16(1) + 16cos(t-2t) + 4(1) = 16 + 16cost + 4 = 20 + 16cost

OpenStudy (anonymous):

I got x^2 + y^2 = 20 + 16cos(t)

OpenStudy (anonymous):

, but now we still have that t , we want to eliminate that

OpenStudy (raden):

yes, twine mine

OpenStudy (anonymous):

twine mine ?

OpenStudy (raden):

nopes... yours like m :)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

where go from here

OpenStudy (raden):

actually, r be scalar... but idk why there is t there :)

OpenStudy (anonymous):

yeah im stumped

OpenStudy (raden):

r^2 = x^2 + y^2 r = sqrt(x^2 + y^2) r = sqrt(20+16cos16t) = sqrt(4(5+4cos16t)) = 2sqrt(5+4cos16t) just simplify it :)

OpenStudy (anonymous):

topic:permutation A building has 5 entrances and 4 exit. In how many ways can you go in and out of the building?

OpenStudy (anonymous):

wait, how did you get sqrt ( 20 + 16 cos 16 ?

OpenStudy (anonymous):

r = sqrt( 20 + 16 cos (theta)) and that is wrong, i just checked the graph

OpenStudy (raden):

pythagorean theorem : r^2 = x^2 + y^2 r = sqrt(x^2 + y^2)

OpenStudy (anonymous):

r = sqrt( 20 + 16 cos (theta)) is not the same as the graph of x = 4sin t + 2sin(2t) , y = 4cos (t) + 2cos(2t)

OpenStudy (raden):

sorry there is typo above :) r = sqrt(20+16cost) = sqrt(4(5+4cost)) = 2sqrt(5+4cost)

OpenStudy (anonymous):

ok, that is the wrong graph, the polar is not the same as the parametric

OpenStudy (anonymous):

I think i got it

OpenStudy (anonymous):

from the second eqution we can solve for cos (t)

OpenStudy (anonymous):

y = 4cos (t) + 2cos(2t) now change cos (2t) = 2 cos^(t) - 1

OpenStudy (anonymous):

so y =4cos(t) + 4 cos^2(t) -2

OpenStudy (anonymous):

then let it equal to zero

OpenStudy (anonymous):

4cos^2(t) + 4cos (t) -2 - y = 0 , solve for cos t

OpenStudy (anonymous):

I got it, but its pretty long solution, message me if youre interested

OpenStudy (raden):

yea, its very interesting... great

OpenStudy (anonymous):

want to see it ?

OpenStudy (raden):

i think u can write it here :)

OpenStudy (anonymous):

so to solve cos t from 4cos^2(t) + 4cos (t) -2 - y = 0 , that

OpenStudy (raden):

yes

OpenStudy (anonymous):

so far we have x^2 + y^2 = 20 + 16 cos(t) and from the second equation we have y = 4cos (t) + 2cos(2t) , substitute cos(2t) = 2cos^2(t) - 1

OpenStudy (anonymous):

now from this 4cos^2(t) + 4cos (t) -2 - y = 0 use quadratic formula, let u = cos t 4u^2 + 4u -2 - y = 0, treat y as a constant , u is the variable u = [-4 + - sqrt( 4^2 - 4*4*[-2-y] ) ] / (2*4)

OpenStudy (raden):

ahhh, yeah i think i have conected ... quadratic formula, i was forget it

OpenStudy (anonymous):

now simplify that

OpenStudy (anonymous):

u = -4/ 8 + - sqrt( 16 + 16(2+y)) / 8

OpenStudy (anonymous):

u = -1/2 + - 4 sqrt(1 + 2+ y ) / 8 u = -1/2 + - 4 sqrt(3+y)/8

OpenStudy (anonymous):

so cos t = -1/2 + - sqrt(3+y)/2

OpenStudy (raden):

so far, i was so good :)

OpenStudy (anonymous):

now lets plug this into x^2 + y^2 = 20 + 16 cos (t)

OpenStudy (anonymous):

x^2 + y^2 = 20 + 16 [-1/2 + - sqrt(3+y)/2 ]

OpenStudy (raden):

right, n then just take sqrt ?, im not sure

OpenStudy (anonymous):

, then distribute

OpenStudy (anonymous):

x^2 + y^2 = 20 + 16*-1/2 +- 16/2 sqrt(3+y)

OpenStudy (anonymous):

x^2 + y^2 = 20 - 8 + - 8 sqrt(3+y)

OpenStudy (anonymous):

x^2 + y^2 = 12 + - 8sqrt(3+y) , now plug in x^2 + y^2 = r^2, and y = r sin theta

OpenStudy (anonymous):

r^2 = 12 + - 8 sqrt(3 + r sin theta) , and thats a polar equation

OpenStudy (anonymous):

and i graphed that on wolfram, and it works. notice that this is is not a function r = f(theta), unless you went real crazy and kept going, r^2 - 12 = + - 8 sqrt(3 + r sin theta) , square both sides

OpenStudy (raden):

make sense for m now, eddie.. what grade ur class, eddie

OpenStudy (anonymous):

im in college

OpenStudy (anonymous):

sophomore

OpenStudy (anonymous):

so honestly i dont think the teacher gave this problem, or if he did , this is very difficult. its a rotated cardiod

OpenStudy (anonymous):

i bet some rotation transformation might make it easier

OpenStudy (raden):

yea, therefore i have to give medal for u today, eddie :) great, that's amazing

OpenStudy (anonymous):

so we learned a lot today i think , never give up . but some problems are intractable , or you need some other method

OpenStudy (anonymous):

new question

OpenStudy (anonymous):

Write the polar equation r = cos(4*theta) in the parametric form x = x(t), y = y(t)

OpenStudy (raden):

i was too old to more studying again...

OpenStudy (anonymous):

we must study further

OpenStudy (anonymous):

where are you from rad?

OpenStudy (raden):

i from indonesia

OpenStudy (raden):

sorry, if my english soo bad :)

OpenStudy (anonymous):

thats ok , so lets multiply both sides by r r = cos (4theta) r^2 = r cos (4theta), but we have a problem, r cos theta = x , not r cos (4theta)

OpenStudy (raden):

sorry, eddie... i want going to my student's home there is a private lesson now, next time we can discuss again

OpenStudy (raden):

nice to know u eddie

OpenStudy (raden):

try post it in new posting, the other members maybe can help u